# A Novel Federated Learning-Based IDS for Enhancing UAVs Privacy and Security

### v2

@@ V-A Classifier: DNN vs CNN
TABLE VI: Accuracy of DNN and CNN in Detecting Blackhole Attack
@@
TABLE VII: Accuracy of DNN and CNN on Detecting Flooding Attack
@@
V-B Effectiveness: Central, Local vs Federated IDS
In this subsection, we compare three intrusion detection approaches: Central IDS (C-IDS), Local IDS (L-IDS), and Federated IDS (FL-IDS) for each attack type. We first discuss the results separately for each attack type and then provide a general discussion.
V-B1 Sinkhole Attack
Three IDSs (C-IDS, L-IDS, and FL-IDS) are compared in Figure 2. At a low attacker ratio of 5%, all IDSs perform poorly. However, their performance significantly improves as the attacker ratio increases. While there is a substantial gap between C-IDS (82.85%) and the others (L-IDS: 72.99%, FL-IDS: 70.44%) at the attacker ratio of 5%, this gap narrows as the attacker ratio in the network increases. Especially at an attacker ratio of 15 or higher, the performance gap between C-IDS and FL-IDS narrows to less than 1.5%. On the other hand, L-IDS consistently performs much worse than the others, even achieves 90.27% accuracy at the highest attacker ratio, while FL-IDS and C-IDS obtain 97.70% and 99.03% accuracy respectively. While some local monitoring nodes such as closely located to the sinkhole in the network might be more affected from the attack, other nodes can be less affected, or even isolated nodes might not be affected at all, hence this situation is reflected in the local IDS’ results.
Figure 2: Comparison of C-IDS, L-IDS and FL-IDS in Detecting Sinkhole Attack
These results highlight the central model’s advantage in accessing data from all clients, enabling high accuracy in detecting sinkhole attacks. However, as mentioned earlier, FL-IDS’s performance converges to a level close to that of C-IDS, with a difference of up to 1.5% at higher attacker ratios ($\geq$ 15%). Figure 3 displays the convergence graphs of each model for attacker ratios of 20% and 25%. The graphs reveal that FL-IDS converges to Central IDS at approximately the 60th epoch.
@@ V-B1 Sinkhole Attack
We also employed an approach called Bias towards specific clients (BTSC) in order to improve the performance of FL-IDS at low attacker ratios. The approach involves giving more weights to the better detecting clients, as some clients may be better at detecting attacks than others due to the dynamic nature of FANETs. In the federated aggregation, 5% of the clients are taken, allowing us to focus on the best detectors. The results obtained with BTSC are shown and compared with the FL-IDS trained with all clients in Table VIII. As shown in the results, when the federated model is trained with the best clients, FL-IDS outperform C-IDS even at the lowest attacker ratio.
TABLE VIII: Performance Improvement with BTSC in Detecting Sinkhole Attack
@@
V-B2 Blackhole Attack
All results for the three IDSs on detecting blackhole attacks are presented in Figure 4. At a low attacker ratio of 5%, all IDSs exhibit inadequate performance in detecting blackhole attacks. The federated learning-based approach, in particular, performs the poorest with a success rate of only 65.03% at this attacker ratio. FL-IDS, being based on decentralized data sources, can only share their weights with the global server. Since attacker nodes cannot influence all nodes uniformly across the network, the data distribution is uneven, leading to lower accuracy at lower attacker ratios. The best performance at the lowest attacker ratio is achieved by C-IDS, reaching 83.68%. C-IDS benefits from its access to all network-generated data, providing a comprehensive view of network traffic and patterns, which contributes to its high accuracy in detecting attacks.
Figure 4: Comparison of C-IDS, L-IDS and FL-IDS in Detecting Blackhole Attack
The performance gap between FL-IDS and C-IDS is diminished as the attacker ratio increased in the network. Specifically, when the attacker ratio reached 25%, the difference in accuracy between C-IDS and FL-IDS reduced to less than 0.5%. Conversely, L-IDS consistently underperformed the other two. The key distinction is that Local IDSs exclusively process their own data and lack a framework for inter-client communication. This absence of communication capabilities accounts for their comparatively lower accuracy.
However, with an increase in the attacker ratio, accuracy tends to improve across all approaches. This is attributed to a higher number of attacker nodes in the network introducing more anomaly data and patterns for deep learning algorithms to analyze. Consequently, all three approaches exhibit increased performance as the attacker ratio rises. As a result, the best performance for L-IDS (90.35%) is achieved at the highest attacker ratio, as expected.
In Figure 5, the accuracy of the IDSs is depicted for each epoch, revealing that FL-IDS closely approached the accuracy of C-IDS at different epochs for various attacker ratios. Notably, FL-IDS reached a level of accuracy comparable to C-IDS at approximately the 30th epoch for the 20% and 25% attacker ratios. However, for the 5% attacker ratio, FL-IDS cannot match C-IDS’s accuracy at any epoch, with C-IDS consistently outperforming FL-IDS in subsequent epochs. Additionally, it is worth noting that L-IDS consistently yielded lower results compared to both types of IDS.
@@ V-B2 Blackhole Attack
The performance of FL-IDS is enhanced by employing the BTSC method as shown in Table IX. When FL-IDS is trained with the weight of the best clients, it achieves high accuracy results, surpassing C-IDS. On the other hand, while its performance reaches 80% accuracy at the lowest ratio, C-IDS still achieves the best performance.
TABLE IX: Performance Improvement with BTSC in Detecting Blackhole Attack
@@
V-B3 Flooding Attack
Figure 6 shows the comparative result of three IDS approaches for detecting flooding attack. Differently from other attacks IDSs based on central or federated learning-based approaches obtain high accuracy even when the attacker ratio is as low as 10%. C-IDS (98.13%) and FL-IDS (97.78%) have high performance since many nodes are affected by the attack and participate in training. On the other hand, L-IDS (87.41%) show considerably lower performance than C-IDS and FL-IDS at the 10% of attacker ratio. This can be resulted from some nodes which might be less affected by the attack than others due to their positions. For instance, some nodes might receive fewer flooding packets due to being more isolated and having a few neighbour nodes only. However, when the attacker ratio is 15% or higher, L-IDS almost reaches the results of C-IDS and FL-IDS, as the number of nodes affected by the attack also increases.
Figure 6: Comparison of C-IDS, L-IDS and FL-IDS in Detecting Flooding Attack
The results reveal that C-IDS benefits from having access to all clients’ data, enabling it to achieve high accuracy, even at lower attacker ratios. In comparison to other attack scenarios, L-IDS maintains competitive performance to C-IDS and FL-IDS, considering the characteristics of flooding attacks that can affect most of the nodes in the network. Its performance is consistent, and its accuracy values are close to those of other IDSs. Figure 7 shows that L-IDS converges to C-IDS and FL-IDS at early epochs (approximately at the 6th epoch for 20% attacker ratio and at the 11th epoch for 25% attacker ratio). The obtained results demonstrate that all three IDSs show high results and reduce the performance gap among them at earlier epochs.
@@ V-B3 Flooding Attack
Finally, although the accuracy results are already high for flooding attacks, FL-IDS has been further improved with the BTSC method, as indicated in Table X.
TABLE X: Performance Improvement with BTSC in Detecting Flooding Attack
@@
V-B4 Detection Rate vs. False Positive Rate
Table XI provides information about the Detection Rate (DR) and False Positive Rate (FPR) for each type of attack. The results reveal that all IDS models exhibit relatively high false positive rates at the lowest attacker ratio. However, as the attacker ratio increases, the FPR tends to decrease to reasonable levels. Notably, when the network contains a significant number of attackers (around 20% to 25%), each IDS models’ performance approaches an ideal scenario. In flooding attacks, IDS converge toward the ideal scenario even when the attacker ratio is as low as 10% in the network, since this attack causes significant harm to the network. Unlike other attacks, flooding attacks affect nearly all nodes, making it easier to differentiate abnormal data from normal data. It is worth highlighting that, on average, blackhole attacks perform slightly better than sinkhole attacks. While these attacks share similarities in directing traffic through attacker nodes, blackhole attacks also drop the packets they attract. This action might lead to an interruption in network traffic, as legitimate packets fail to reach their intended destinations. This disruption in network communications is more readily detectable because it directly impacts connectivity and traffic flow.
TABLE XI: Evaluation Results of C-IDS, L-IDS and FL-IDS in terms of DR and FPR
@@ VI-A Accuracy
The results discussed above clearly show that when we centrally obtain local data and train a model using all these data, we can better differentiate attacks from benign network traffic. Therefore, Central IDS (C-IDS) demonstrates the best performance for each attack type. On the other hand, Federated IDS (FL-IDS) achieves competitive results with C-IDS. Particularly at higher attacker ratios, FL-IDS can approach the accuracy of C-IDS. Because, when the attacker ratio increases, more devices affected by such attacks can positively participate in training of FL-IDS. This is also supported by the improvement obtained with the use of BTSC, which selectively considers the best clients for training FL-IDS.
@@
Local IDSs (L-IDS) have a lower success rate on average compared to other IDSs. The performance of L-IDS is calculated as the average of each local client. Therefore, L-IDS only show comparable results with C-IDS and FL-IDS in detecting flooding attacks due to the nature of this attack type, which affects most of the nodes in the network. However, when the impact of an attack is limited, as is the case with sinkhole and blackhole attacks that affect specific victims and its surroundings, the average accuracy of local IDSs might not be high. This discrepancy arises because while some clients can effectively detect attacks due to their location (such as being situated in active routes or denser areas), others may have less success.
@@
The performance of individual local IDS has been compared in Tables XII, XIII, and XIV for sinkhole, blackhole, and flooding attacks, respectively. These tables display the performance of the best and worst local IDS for both L-IDS and FL-IDS, as well as their average performance. One notable observation is that local ID systems, which receive feedback from the global model as part of federated learning, exhibit significantly superior performance compared to traditional L-IDS. FL-IDS effectively leverages the advantages of both centralized and local approaches, resulting in competitive accuracy levels across a variety of attacker ratios. However, it is worth noting that L-IDS outperforms FL-IDS at the lowest attacker ratios for blackhole and sinkhole attacks. In contrast, when dealing with flooding attacks and as the number of affected nodes in the network increases, FL-IDS shows better performance than L-IDS even at the lowest attacker ratio. Hence, it is recommended to employ BTSC for training IDS in the context of federated learning. This approach results in improved performance, particularly when a significant portion of the network’s clients remains unaffected by the attack.
@@
TABLE XII: Best/Worst Performance of Local IDSs in Detecting in Sinkhole Attacks
@@
TABLE XIII: Best/Worst Performance of Local IDSs in Detecting in Blackhole Attacks
@@
TABLE XIV: Best/Worst Performance of Local IDSs in Detecting in Flooding Attacks
@@
To summarize, FL-IDS demonstrates high performance in detecting a range of network attacks, including blackhole, sinkhole, and flooding attacks. Notably, when it is trained with the BTSC, FL-IDS even approaches the performance level of traditional Centralized IDS (C-IDS). This is achieved without the need for centrally collecting all local data, showcasing the remarkable effectiveness of federated learning in network security. Through FL-IDS, local clients can learn and adapt independently while benefiting from global model updates. In addition to improving overall performance, this collaboration helps in addressing complex and evolving threats that individual clients might not have encountered. With periodic retraining and ongoing evolution of the global model, FL-IDS ensures that it stays current and effective in countering both existing and emerging threats, while seamlessly welcoming new nodes into the network.
VI-B Communication and Response

### v3

@@ 5.1 Classifier: DNN vs CNN
Table 6: Accuracy of DNN and CNN in Detecting Blackhole Attack
@@
Table 7: Accuracy of DNN and CNN on Detecting Flooding Attack
@@
5.2 Effectiveness: Central, Local vs Federated IDS
In this subsection, we compare three intrusion detection approaches: Central IDS (C-IDS), local IDS (L-IDS), and federated IDS (FL-IDS) for each type of attack. We first discuss the results separately for each attack type and then provide a general discussion.
5.2.1 Sinkhole Attack
Three IDSs (C-IDS, L-IDS, and FL-IDS) are compared in Figure 3(a). With a low attacker ratio of 5%, all IDSs perform poorly. However, their performance improves significantly as the attacker ratio increases. Although there is a substantial gap between C-IDS (82.85%) and the others (L-IDS: 72. 99%, FL-IDS: 70. 44%) in the attacker ratio of 5%, this gap narrows as the attacker ratio in the network increases. Especially at an attacker ratio of 15 or higher, the performance gap between C-IDS and FL-IDS narrows to less than 1.5%. In federated learning, a higher proportion of attacker nodes exposes individual nodes to a wider range of anomalous data and attack patterns, increasing their exposure to attack behaviors. This broader exposure improves the detection accuracy at individual nodes, allowing them to learn and adapt to a more diverse set of attack characteristics. Consequently, the global model, which aggregates these enhanced local models, demonstrates improved overall effectiveness in detecting attacks, thereby reducing the performance difference between FL-IDS and C-IDS as attacker ratios increase. On the other hand, L-IDS consistently performs much worse than the others, even achieves 90.27% accuracy at the highest attacker ratio, while FL-IDS and C-IDS obtain 97. 70% and 99. 03% accuracy, respectively. While some local monitoring nodes such as close to the sinkhole in the network might be more affected from the attack, other nodes can be less affected, or even isolated nodes might not be affected at all, hence this situation is reflected in the local IDS’ results.
(a) Sinkhole Attack
(b) Blackhole Attack
(c) Flooding Attack
Figure 3: Comparison of C-IDS, L-IDS, and FL-IDS in Detecting Attacks. Subfigures (a), (b), and (c) show individual attack types.
@@ 5.2.1 Sinkhole Attack
These results highlight the central model’s advantage in accessing data from all clients, enabling high accuracy in detecting sinkhole attacks. However, as mentioned earlier, the performance of FL-IDS converges to a level close to that of C-IDS, with a difference of up to 1.5% at higher attacker ratios ($\geq$ 15%). Figure 4 displays the convergence graphs of each model for attacker ratios of 20% and 25%. The graphs reveal that FL-IDS converges to Central IDS at approximately the 60th epoch.
We also employed an approach called Bias towards specific clients (BTSC) in order to improve the performance of FL-IDS at low attacker ratios. The approach involves giving more weights to the better detecting clients, as some clients may be better at detecting attacks than others due to the dynamic nature of FANETs. In the federated aggregation, 20% of the clients are taken, allowing us to focus on the best detectors. The results obtained with BTSC are shown and compared with the FL-IDS trained with all clients in Table 8. As shown in the results, when the federated model is trained with the best clients, FL-IDS outperform C-IDS even at the lowest attacker ratio.
Table 8: Performance Improvement with BTSC in Detecting Sinkhole Attack
@@
5.2.2 Blackhole Attack
All results for the three IDSs on the detection of blackhole attacks are presented in Figure 3(b). With a low attacker ratio of 5%, all IDSs exhibit inadequate performance in detecting blackhole attacks. The federated learning-based approach, in particular, performs the worst with a success rate of only 65.03% at this attacker ratio. FL-IDS, being based on decentralized data sources, can only share their weights with the global server. Since attacker nodes cannot influence all nodes uniformly across the network, the data distribution is uneven, leading to lower accuracy at lower attacker ratios. The best performance at the lowest attacker ratio is achieved by C-IDS, reaching 83.68%. C-IDS benefits from its access to all network-generated data, providing a comprehensive view of network traffic and patterns, which contributes to its high accuracy in detecting attacks.
The performance gap between FL-IDS and C-IDS decreases as the attacker ratio increases in the network. Specifically, when the attacker ratio reached 25%, the difference in accuracy between C-IDS and FL-IDS reduced to less than 0.5%. In contrast, L-IDS consistently underperformed the other two. The key distinction is that Local IDSs exclusively process their own data and lack a framework for inter-client communication. This lack of communication capabilities accounts for their comparatively lower accuracy.
However, with an increase in the attacker ratio, accuracy tends to improve across all approaches. This is attributed to a higher number of attacker nodes in the network, introducing more anomaly data and patterns for deep learning algorithms to analyze. Consequently, all three approaches exhibit increased performance as the attacker ratio increases. As a result, the best performance for L-IDS (90. 35%) is achieved with the highest attacker ratio, as expected.
In Figure 5, the accuracy of the IDSs is depicted for each epoch, revealing that FL-IDS closely approached the accuracy of C-IDS at different epochs for various attacker ratios. In particular, FL-IDS reached a level of accuracy comparable to C-IDS at approximately the 30th epoch for the 20% and 25% attacker ratios. However, for the 5% attacker ratio, FL-IDS cannot match C-IDS’s accuracy at any epoch, with C-IDS consistently outperforming FL-IDS in subsequent epochs. Furthermore, it should be noted that L-IDS consistently yielded lower results compared to both types of IDS.
@@ 5.2.2 Blackhole Attack
The performance of FL-IDS is improved by employing the BTSC method as shown in Table 9. When FL-IDS is trained with the weight of the best clients, it achieves high accuracy results, surpassing C-IDS. On the other hand, while its performance reaches 80% accuracy at the lowest ratio, C-IDS still achieves the best performance.
Table 9: Performance Improvement with BTSC in Detecting Blackhole Attack
@@
5.2.3 Flooding Attack
Figure 3(c) shows the comparative result of three IDS approaches for detecting flooding attack. Differently from other attacks IDSs based on central or federated learning-based approaches obtain high accuracy even when the attacker ratio is as low as 10%. C-IDS (98. 13%) and FL-IDS (97. 78%) have a high performance since many nodes are affected by the attack and participate in training. On the other hand, L-IDS (87.41%) show considerably lower performance than C-IDS and FL-IDS in the 10% of the attacker ratio. This can be a result of some nodes which might be less affected by the attack than others due to their positions. For instance, some nodes might receive fewer flooding packets due to being more isolated and having a few neighbour nodes only. However, when the attacker ratio is 15% or higher, L-IDS almost reaches the results of C-IDS and FL-IDS, as the number of nodes affected by the attack also increases.
The results reveal that C-IDS benefits from having access to all client data, enabling it to achieve high accuracy, even at lower attacker ratios. In comparison to other attack scenarios, L-IDS maintains competitive performance to C-IDS and FL-IDS, considering the characteristics of flooding attacks that can affect most of the nodes in the network. Its performance is consistent, and its accuracy values are close to those of other IDSs. Figure 6 shows that L-IDS converges to C-IDS and FL-IDS at early epochs (approximately at the 6th epoch for 20% attacker ratio and at the 11th epoch for 25% attacker ratio). The obtained results demonstrate that all three IDSs show high results and reduce the performance gap among them at earlier epochs.
@@ 5.2.3 Flooding Attack
Finally, although the accuracy results are already high for flooding attacks, FL-IDS has been further improved with the BTSC method, as indicated in Table 10.
Table 10: Performance Improvement with BTSC in Detecting Flooding Attack
@@
5.2.4 Detection Rate vs. False Positive Rate
Table 11 provides information on the detection rate (DR) and the false positive rate (FPR) for each type of attack. The results reveal that all IDS models exhibit relatively high false positive rates at the lowest attacker ratio. However, as the attacker ratio increases, the FPR tends to decrease to reasonable levels. In particular, when the network contains a significant number of attackers (around 20% to 25%), each IDS model’s performance approaches an ideal scenario. In flooding attacks, IDS converge to the ideal scenario even when the attacker ratio is as low as 10% on the network, since this attack causes significant damage to the network. Unlike other attacks, flooding attacks affect almost all nodes, making it easier to differentiate abnormal data from normal data. It is worth highlighting that, on average, blackhole attacks perform slightly better than sinkhole attacks. While these attacks share similarities in directing traffic through attacker nodes, blackhole attacks also drop the packets they attract. This action might lead to an interruption in network traffic, as legitimate packets fail to reach their intended destinations. This disruption in network communications is more readily detectable because it directly impacts connectivity and traffic flow.
Table 11: Evaluation Results of C-IDS, L-IDS and FL-IDS in terms of DR and FPR
@@ 6.1 Accuracy
The results discussed above clearly show that when we centrally obtain local data and train a model using all these data, we can better differentiate attacks from benign network traffic. Therefore, Central IDS (C-IDS) shows the best performance for each type of attack. However, Federated IDS (FL-IDS) achieves competitive results with C-IDS. In particular, at higher attacker ratios, FL-IDS can approach the accuracy of C-IDS because the increased number of attacker nodes introduces a richer set of anomalous patterns across the network. This leads to more robust local models being trained in the data set, which are then aggregated into a global model that better generalizes attack detection across the entire network. When the attacker ratio increases, more devices affected by such attacks can positively participate in FL-IDS training. This is also supported by the improvement obtained with the use of BTSC, which selectively considers the best clients for training FL-IDS.
@@
Local IDSs (L-IDS) have a lower success rate on average compared to other IDSs. The performance of L-IDS is calculated as the average of each local client. Therefore, L-IDS only shows results comparable to those of C-IDS and FL-IDS in detecting flooding attacks due to the nature of this attack type, which affects most nodes in the network. However, when the impact of an attack is limited, as is the case with sinkhole and blackhole attacks that affect specific victims and their surroundings, the average accuracy of local IDSs may not be high. This discrepancy arises because, while some clients can effectively detect attacks due to their location (such as being situated in active routes or denser areas), others may have less success.
The performance of IDS on the nodes has been compared in Tables 12, 13, and 14 for sinkhole, blackhole, and flooding attacks, respectively, based on accuracy. These tables present the performance of the best and worst IDSs for both L-IDS and FL-IDS when deployed on different nodes, along with their average performance. In traditional L-IDS, where each node operates solely on its local data, we observe a significant disparity between the worst and best node performance of detection. This variation underscores the limitations of L-IDS, particularly in scenarios where nodes have insufficient data or lack diversity in their training sets. FL-IDS outperforms L-IDS in terms of the worst-performing node across all attacker ratios and attack types. However, the best-performing node in L-IDS achieves a higher detection accuracy than the best-performing node in FL-IDS. This difference can be attributed to the fact that in L-IDS, each node is trained exclusively on its own local data. As a result, when the test data closely matches the node’s local attack patterns, L-IDS can achieve significantly higher detection performance. However, this advantage is node specific and highly dependent on the distribution and density of local attack data.
@@
One notable observation is that under blackhole and sinkhole attacks, when the attacker ratio exceeds 10%, worst-performing node in FL-IDS that receive feedback from the global model as part of federated learning demonstrate better performance than the average accuracy of L-IDS. However, when dealing with flooding attacks and as the number of affected nodes increases, L-IDS outperforms FL-IDS, even at lower attacker ratios. To further enhance the performance of FL-IDS, it is recommended to use BTSC for training IDS in federated learning, particilarly in scenerios where a significant portion of the network’s clients remain unaffected by the attack.
The use of BTSC results in improved performance, particularly when a significant portion of the network’s clients remain unaffected by the attack. By prioritizing high-performing clients based on their detection accuracy, BTSC ensures that only the most relevant and generalizable model updates are incorporated during the federated aggregation process. This leads to significant improvements in detection accuracy. In addition, BTSC reduces communication overhead by limiting the number of contributing clients, which is particularly beneficial in resource-constrained FANET environments. Additionally, the approach helps lower energy consumption by minimizing the need for extensive data transmission. The overall performance improvements, along with the robust generalization demonstrated through validation on a diverse test dataset, underscore the validity of the proposed approach in improving the efficiency and effectiveness of FL-IDS in dynamic and heterogeneous FANETs.
@@
Table 12: Best/Worst Accuracy of IDSs in Detecting in Sinkhole Attacks
@@
Table 13: Best/Worst Accuracy of IDSs in Detecting in Blackhole Attacks
@@
Table 14: Best/Worst Accuracy of IDSs in Detecting in Flooding Attacks
@@
We also evaluate the performance of three widely used aggregation algorithms, FedAvg, FedProx [62], and FedSGD [63] . Our evaluation shows that FedAvg consistently outperforms FedProx across key metrics, including accuracy, DR, and FPR. FedProx is designed to handle environments with significant systems heterogeneity and statistical heterogeneity [62]. In our setting, however, FedProx’s regularization mechanism appears to introduce additional constraints that slow convergence and diminish attack detection performance, leading to worse results compared to FedAvg. FedSGD exhibits distinct performance characteristics compared to both FedAvg and FedProx. Unlike these methods, FedSGD updates the global model after each mini-batch, resulting in increased communication frequency but reduced local computation. The results show that FedSGD maintains an exceptionally low FPR; however, its DR is significantly lower than both FedAvg and FedProx, indicating weaker performance in identifying attacks. This reduced DR may stem from the lack of local model training, limiting the model’s ability to learn robust attack patterns.
Table 15: Comparative Evaluation of Aggregation Algorithms
FedAvg FedProx FedSGD
Attack Type Attack ratio (%) Accuracy DR FPR Accuracy DR FPR Accuracy DR FPR
Sinkhole 5 70.44% 83.00% 17.00% 56.96% 69.00% 31.00% 56.97% 0.54% 0.45%
10 89.18% 95.67% 4.33% 55.85% 49.00% 51.00% 66.56% 0.61% 0.38%
15 94.07% 98.01% 1.99% 65.41% 58.00% 42.00% 71.34% 0.68% 0.31%
20 97.41% 99.31% 0.69% 65.41% 71.00% 29.00% 72.65% 0.79% 0.20%
25 97.70% 99.48% 0.52% 66.74% 69.00% 31.00% 74.65% 0.77% 0.22%
Blackhole 5 65.04% 75.17% 24.83% 58.67% 60.00% 40.00% 52.93% 0.42% 0.57%
10 90.00% 93.10% 6.90% 61.33% 57.00% 43.00% 63.01% 0.54% 0.45%
15 95.10% 95.16% 4.84% 64.75% 39.00% 61.00% 64.44% 0.47% 0.52%
20 99.04% 99.48% 0.52% 66.44% 47.00% 53.00% 74.14% 0.67% 0.32%
25 99.26% 99.66% 0.34% 68.59% 73.00% 27.00% 82.72% 0.85% 0.14%
Flooding 5 76.12% 63.44% 36.56% 58.43% 45.00% 55.00% 66.21% 0.59% 0.40%
10 97.78% 98.14% 1.86% 78.67% 62.00% 38.00% 86.76% 0.79% 0.20%
15 98.30% 98.79% 1.21% 80.22% 80.00% 20.00% 86.84% 0.83% 0.16%
20 99.26% 99.48% 0.52% 80.30% 66.00% 34.00% 87.89% 0.85% 0.14%
25 99.33% 99.61% 0.39% 74.74% 74.00% 26.00% 89.37% 0.86% 0.13%
@@
In summary, FL-IDS shows high performance in detecting a range of network attacks, including black-hole, sinkhole, and flood attacks. In particular, when trained with the BTSC, FL-IDS even approaches the performance level of traditional Centralized IDS (C-IDS). This is achieved without the need to centrally collect all local data, showcasing the remarkable effectiveness of federated learning in network security. Through FL-IDS, local clients can learn and adapt independently while benefiting from global model updates. In addition to improving overall performance, this collaboration helps address complex and evolving threats that individual clients may not have encountered. With periodic retraining and an ongoing evolution of the global model, FL-IDS ensures that it stays current and effective in countering both existing and emerging threats, while seamlessly welcoming new nodes into the network.
6.2 Communication and Response